Refraction through glass and water


A plate of glass with parallel faces having a refractive index of 1.58 is resting on the surface of water in a tank. A ray of light coming from above in air makes an angle of incidence 37.5?with the normal to the top surface of the glass.

What angle ?3 does the ray refracted into the water make with the normal to the surface? Use 1.33 for the index of refraction of water.

Express your answer in degrees
Answer #1

N2/N1 = sin(i)/sin(r)

Thus, sin(r) = N1(sin(i))/N2 = 1.58(sin37.5)/1.33 = 0.70136...

So angle(r) = 44.5 degrees